### v1

# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

Abstract
Updates of extensive Internet of Things (IoT) data are critical to the immersion of vehicular metaverse services. However, providing high-quality and sustainable data in unstable and resource-constrained vehicular networks remains a significant challenge. To address this problem, we put forth a novel immersion-aware model trading framework that incentivizes metaverse users (MUs) to contribute learning models trained by their latest local data for augmented reality (AR) services in the vehicular metaverse, while preserving their privacy through federated learning. To comprehensively evaluate the contribution of locally trained learning models provided by MUs to AR services, we design a new immersion metric that captures service immersion by considering the freshness and accuracy of learning models, as well as the amount and potential value of raw data used for training. We model the trading interactions between metaverse service providers (MSPs) and MUs as an equilibrium problem with equilibrium constraints (EPEC) to analyze and balance their costs and gains. Moreover, considering dynamic network conditions and privacy concerns, we formulate the reward decisions of MSPs as a multi-agent Markov decision process. Then, a fully distributed dynamic reward method based on deep reinforcement learning is presented, which operates without any private information about MUs and other MSPs. Experimental results demonstrate that the proposed framework can effectively provide higher-value models for object detection and classification in AR services on real AR-related vehicle datasets compared to benchmark schemes.
Index Terms:
Equilibrium problem with equilibrium constraints, immersion-aware, incentive mechanism, resource allocation, vehicular metaverse.
I Introduction
The metaverse, envisioned as one of the next evolution directions of the Internet, is designed to create a fully immersive, self-sustaining virtual environment in which people can engage in activities such as playing, working, and socializing [1]. This vision is propelled by advances in the fifth-generation (5G) and the forthcoming sixth-generation (6G) communication technologies, renowned for their low latency and high data throughput. These technologies play a critical role in seamlessly integrating the Internet of Things (IoT) data into metaverse services, thus bringing the once-fictional concept of immersive experiences closer to reality.
Metaverse services are beginning to reveal their vast potential across a broad spectrum of industries, from gaming and autonomous driving to education and marketing. Notably, the application of vehicles within the metaverse has captured significant interest due to the enhanced safety features and immersive experiences offered by state-of-the-art augmented reality (AR) technologies.
Figure 1: An example of AR services in the vehicular metaverse [^2] Image source: https://www.jasoren.com/ar-in-automotive/.
Market report [2] forecasted that the global automotive metaverse market will grow from $1.9$ billion in 2022 to $6.5$ billion by 2030. Automakers like BMW have recently doubled down on AR technology in their projects. As shown in Fig. 2, augmented information makes driving safer by showing potential hazards hidden behind the vehicle in front of you. Moreover, Nissan’s [3] upcoming technology utilizes a 3D AR interface that merges the real world with the virtual world to provide the driver with augmented information about the surrounding area. In addition, Vanarama has also implemented a visual parking assistance function using AR [4], which is expected to identify empty spaces in busy car parks and display useful information, such as parking fees, on the windscreen.
The ability to capture information from the “real” world, particularly the ability to collect and process massive data from IoT devices, is the key to determining the success of immersive services (e.g., AR) in the vehicular metaverse. Meanwhile, the data must be processed and presented in a meaningful, responsive, and appropriately protected way. Technically, a high-quality experience with AR services relies on accurate detection and classification of real-world objects (e.g., cars and pedestrians) under complex conditions [5]. To achieve this goal, sufficient valid data needs to be collected and processed in depth to detect and classify objects accurately. Therefore, it is essential to focus on effectively collecting, processing, and protecting the data that supports a safer and more enjoyable driving experience.
Motivations. An existing widely-used data collection method, as adopted by Nissan [6], involves collecting massive amounts of data through vehicle sensors, cameras, and roadside devices, and then processing all the data centrally. However, when it comes to the situation with multiple metaverse users (MUs) and metaverse service providers (MSPs) on behalf of different companies, the centralized data collection approach is not applicable and may lead to the following issues. First, AR services in the vehicular metaverse need to be highly immersive so that MUs feel fully immersed in the rendered environment, such as visualized driving. However, providing a seamless AR experience is challenging due to the latency caused by massive real-time data updates under unstable and resource-limited communication conditions. Note that the value of real-time data diminishes over time [7]. Also, delays can severely impact the MU’s experience and cause dizziness [8]. Second, the data to be transmitted may be sensitive and private, such as location, movement, and biometrics, which can create a better immersive experience but may inevitably increase the privacy risk of MUs [9]. Third, it is difficult to share data between MSPs due to competition. This leads to the same data being sampled multiple times by different MSPs in the physical world, resulting in wasted resources (e.g., sampling costs) and data redundancy. Moreover, tremendous data pressure and computational burden are imposed on MSPs. Therefore, an efficient method for collecting high-quality data while preserving privacy is needed to achieve immersive AR services.
Challenges. To address the above issues, we propose an immersion-aware model trading framework designed to provide data support for AR services and ensure privacy by integrating federated learning (FL). Adopting FL enables MUs to contribute diverse learning models to multiple MSPs using local data, thus effectively protecting sensitive information [10]. Moreover, learning models are uploaded to MSPs for AR services by individual MUs, rather than centrally transmitting large amounts of data, which significantly reduces the communication burden. Nonetheless, developing such a model trading framework still faces the following challenges:
  • MUs are typically self-interested and are reluctant to share learning models with MSPs due to the additional computation, communication, and energy overhead. Thus, the challenge here is how to incentivize them to become contributors in providing high-value learning models that will benefit vehicular metaverse services.
  • It is difficult to design an appropriate metric to comprehensively evaluate the value of learning models contributed by MUs due to the diverse aspects involved in model valuation.
  • MUs have different sampling costs and limited computation and communication resources, while MSPs differ in their model preferences and compete with each other. Hence, the next challenge is to model the dynamic competitive interactions among MSPs and to achieve an equilibrium of interests among MUs and MSPs in the model trading.
Figure 2: The outline of an immersion-aware framework for FL-assisted vehicular metaverse.
Contributions. To address these challenges, we model the dynamic competitive trading interactions as an equilibrium problem with equilibrium constraints (EPEC) involving multiple leaders (MSPs) and multiple followers (MUs), which is a hierarchical optimization problem with equilibria at two levels [11]. Furthermore, to evaluate the value of learning models contributed by MUs to MSPs, we design a new metric called the immersion of the learning model (IoM). The metric jointly considers the freshness and accuracy of the learning model, as well as the amount and potential value of raw data used for training. Moreover, given the dynamic networks and the privacy concerns of MSPs, we formulate the reward decisions of MSPs as a multi-agent Markov decision process (MAMDP) and develop a multi-agent deep reinforcement learning (DRL)-based dynamic reward (MDDR) approach to obtain the reward decisions in a fully distributed manner. The outline of the proposed framework is shown in Fig. 2, and the main contributions of this paper are summarized as follows:
  • We propose an immersion-aware trading framework for learning models with FL assistance, incentivizing MUs to contribute high-value learning models for MSPs in a privacy-preserving and efficient manner. Moreover, a novel metric called “IoM” is designed to quantify the immersion enhancement provided by MUs for AR services.
  • We model the dynamic competitive trading interactions as an EPEC and provide theoretical support for the existence and uniqueness of equilibria at two levels. Moreover, a fully distributed MDDR approach is developed to adapt to dynamic environments and address complex reward decisions of MSPs without accessing any private information of MUs or MSPs.
  • We conduct extensive numerical simulations based on AR-related vehicle datasets to validate the efficacy and efficiency of MDDR and the proposed immersion-aware model trading framework. Compared to benchmark schemes, the framework better motivates MUs to provide higher-value learning models for object detection and classification functions of AR services effectively.
The rest of this paper is organized as follows. Section II discusses the related work. In Section III, we present the system overview and design the immersion metric of the learning model. Section IV gives the game formulation, and Section V analyzes the existence of the equilibria at two levels. In Section VI, we give the detailed design of MDDR. Section VII shows numerical experiments to evaluate the framework performance, and finally Section VIII concludes the paper.
II Related Work
Incentive Mechanism. Incentive Mechanisms have been widely applied to metaverse services [12, 13, 14, 15, 16, 17]. For example, Xu et al. [12] proposed an incentive framework for learning-based virtual reality (VR) services and designed a perceptual quality to measure VR users’ immersion in the virtual world. Sun et al. [13] investigated dynamic digital twins in air-assisted vehicular networking and designed a two-stage resource allocation incentive mechanism based on the Stackelberg game. Jiang et al. [14] developed a hierarchical game-theoretic coded distributed computing (CDC) framework for metaverse services. The framework employs coalition and Stackelberg games to select reliable CDC workers for MSPs and incentivize reliable CDC workers to participate in metaverse services. Xi et al. [16] proposed an incentive-based congestion control scheme for stochastic demand response in digital twin edge networks, using Lyapunov optimization and contract theory. Nguyen et al. [17] introduced MetaChain, an innovative blockchain-based framework designed to address the emerging challenges in developing metaverse applications. The framework integrates game theory to motivate MUs to contribute to the metaverse.
However, the above incentives for metaverse services mainly focus on digital twins, blockchain, or virtual reality, while ignoring the scenarios of multiple MSPs and multiple MUs in the vehicular metaverse. In this scenario, the complexity of interactions among MSPs and MUs requires novel incentive mechanisms to better facilitate the development and innovation of metaverse services.
AR and FL. The AR applications, characterized by smoothness and immersion requirements, provide immersive and interactive experiences for MUs by seamlessly integrating virtual content with the real world. For a better immersion experience, Cozzolino et al. [18] developed a latency-energy efficient task offloading for AR services and designed a best-effort resource provisioning algorithm to solve the problem of heterogeneous resource competition. Further, to reduce latency while protecting data privacy, Chen et al. [19] proposed a framework that integrates mobile edge computing and FL to solve the computational efficiency and low-latency object classification problem for AR services. In FL, MUs can train AI models on their local devices, thus enabling privacy-preserving collaborative machine learning while fully leveraging the computation resources of those MUs [20]. Furthermore, Lu et al. [21] investigated communication-efficient FL for digital twin edge networks in industrial IoT, which provides virtual and real synchronization while protecting privacy. Incorporating FL into AR services brings the benefits of reduced communication latency and privacy protection. However, the existing studies fail to consider the integration of FL into AR for the multi-MSP and multi-MU metaverse scenario, while overlooking the selfishness of MUs, i.e., whether they are willing to participate in FL learning.
Inspired by the above existing works, we design an immersion-aware framework integrated FL to provide a learning model trading platform for MSPs and MUs in the vehicular metaverse service markets, enabling MUs to contribute high-value learning models for AR services.
Figure 3: The workflow of the immersion-aware model trading framework [^4] Image source: https://mixed-news.com/en/google-transforms-google-maps-into-the-backbone-of-our-ar-future/, https://about.facebook.com/meta, https://www.vanarama.com/blog/cars/4-ways-augmented-reality-will-revolutionise-the-automotive-industry..
III System Model
We first present the system overview of the immersion-aware model trading framework in Section III-A. We next introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
III-A System Overview
The workflow of the immersion-aware model trading framework is illustrated in Fig. 4, aiming to provide MUs with immersive and interactive AR service experiences. MUs can contribute learning models to the AR services of MSPs for better immersive experiences. For example, services such as object detection can be improved by training on images of streets, pedestrians, and so on, captured by vehicles’ cameras. In this way, AR services can be better applied to visible driving, with more accurate and timely hazard warnings from windscreens. The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services. MSPs are fueled by many technologies that work together to support vehicular metaverse services, including AI engines, AR/VR engines, metaverse trading engines, etc. The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs. MUs play essential roles as both consumers and contributors to vehicular metaverse services.
TABLE I: Parameter Definitions and Notations
Notation Definition
$M$ Number of MUs
$N$ Number of MSPs
$\overline{\Delta}_{mn}$ The average AoI of MU $m$’s model provided for MSP $n$
$V_{mn}$ The IoM contributed by MU $m$ to MSP $n$
$f_{mn}$ MU $m$’s computation resource used for MSP $n$
$B_{mn}$ MU $m$’s communication resource used for MSP $n$
$X_{mn}$ MU $m$’s training data set for MSP $n$
$\omega_{mn}$ The potential value of MU $m$’s local data for MSP $n$
$\theta_{m}$ MU $m$’s local accuracy threshold
The framework is a multi-engine configuration, and this paper focuses on two components: the trading engine and the AI engine of the MSP. The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computation and communication resource allocation of MUs and reward decisions of MSPs). Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide learning models for MSPs under the strategic guidance of the trading results. The specific workflow of the framework involves the following two phases.
Phase I (Incentive process): Initially, MSPs’ AI engines broadcast their global models to all MUs in the vehicular metaverse [^5] The FL global models owned by different MSPs we consider could be object detection or classification models for intelligent terrain mapping, intuitive road safety, visible driving, etc. Note that it is feasible to train on multiple tasks with the same local data. For example, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes. (① in Fig. 4). Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②). Next, MUs determine their allocation of computation and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs. (③). Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their benefits. Steps ② and ③ are repeated until an agreement is reached among MUs and MSPs.
Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computation resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④). Then, MUs upload their updated learning models to the corresponding AI engines of MSPs by consuming communication resources (⑤). Following this, AI engines aggregate models from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥). Finally, the AR/VR engines enable better AR services with the assistance of the AI engines (⑦). Steps ④-⑦ are repeated until the guidance time $T$ ends. To ease the presentation, we summarize some important notations in Table I.
III-B FL Framework Mechanism Adopted
Here, we consider a set of =$\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs with the capacities of local computation and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^6] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive updated models from all MUs before performing model aggregation. The advantages of synchronous FL are high model accuracy and fast convergence [22, 23].. Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}(n\in\mathcal{N})$. The MUs can join FL to generate local learning models and contribute to enhancing vehicular metaverse services. As shown in Fig. 4, MU $m$ provides learning models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel. Moreover, the data collected at the $(r-1)$-th round is used as the local data for training at the $r$-th round. Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of the period $T$.
Figure 4: An illustration of the FL mechanism with $n$ tasks for MU $m$.
To be concrete, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$. The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles. The steps involved in each iteration $r\in\left\{1,2,\ldots,R\right\}$ are as follows:
  • Publishing Tasks : The AI engine of MSP $n$ broadcasts its global model $\jmath_{n}^{(r)}$ (e.g., object detection models used in AR services for visible driving and intuitive road safety) to all MUs in the $r$-th round.
  • Local Training: Each MU, for example, MU $m$, trains $\jmath_{n}^{(r)}$ with the most recently collected data $X_{mn}^{(r)}$ by stochastic gradient descent (SGD) within certain local rounds. The number of local training rounds depends on its local accuracy threshold.
  • Uploading and Aggregation: MUs transmit their updated local model $\jmath_{mn}^{(r)}$ to the AI engine of MSP $n$. After receiving all the updated models, the MSP $n$’s AI engine aggregates them through the following weighted average:
    $\jmath_{n}^{(r+1)}$ $=\jmath^{(r)}_{n}+\sum_{m=1}^{M}\frac{|X_{mn}^{(r)}|}{|X_{n}^{(r)}|}(\jmath_{n}^{(r)}-\jmath_{mn}^{(r)}),$ (1)
    and then obtains the new global model $\jmath_{n}^{(r+1)}$ for the next iteration.
  • Model Utilization: After global aggregation, MSP $n$’s AR engine obtains an updated global model that can be used for corresponding 3D modeling and other functions [19]. The global rounds are iterated until a specific requirement is met, such as reaching a certain level of accuracy or a deadline.
III-C Immersion Metric of Learning Model
MSPs determine the rewards for MUs based on the immersion of the learning model (IoM) metric. Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computation and communication resources to maximize the benefits. This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making. Unlike traditional FL and incentive mechanisms, in vehicular metaverse scenarios with immersion requirements, both the contribution and freshness of the learning model need to be considered [^7] For example, an MSP may receive fresher learning models but with lower contributions, or less fresh learning models but with higher contributions. , which will collectively affect the immersive experience of AR services, i.e., whether virtual objects can be placed in the physical world accurately and promptly. With the above consideration, we propose the IoM, a metric that jointly considers the contribution prediction $I_{mn}$ and the age of information (AoI) $\Delta_{mn}$ [24] to measure the value of a learning model that MU $m$ brought to the vehicular metaverse service of MSP $n$, denoted as
$V_{mn}=I_{mn}(\tau_{n}-\overline{\Delta}_{mn}),$ (2)
where the components are detailed as follows. In the rest of this paper, we refer to the trading engine of MSP as MSP.
III-C1 Contribution Prediction $I$
The contribution prediction $I_{mn}$ from MU $m$ to MSP $n$ is determined by the accuracy $\theta_{m}$ of the learning model, the total amount of training data $\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|$, and the potential value $\omega_{mn}$ of local data. Prior work [25] characterized the contribution of a learning model in terms of a logarithmic function of the amount of local training data. However, evaluating the contribution solely by the training data size is one-sided, as there may be a large amount of redundant data or the validity of the learning model cannot be guaranteed. It is more practical to comprehensively evaluate the amount of training data, model accuracy, and the potential value of local data; thus, we denote $I_{mn}$ as
$I_{mn}=\frac{\omega_{mn}\epsilon ln(1+\eta\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|)}{\theta_{m}},$ (3)
where $\lfloor\frac{T}{\tau_{n}}\rfloor$ is the number of iterations that the task of MSP $n$ can be performed within $T$. $\epsilon$ and $\eta$ are system parameters from experiments as obtained in [26]. $\theta_{m}\in\left(0,1\right)$ characterizes the accuracy of local training as decided by the MU. Here, $\theta\rightarrow 0$ means that a high-accuracy learning model is available, while $\theta\rightarrow 1$ implies that the accuracy of the learning model is low. The local data potential value $\omega_{mn}$ describes the difference between model predictions and true labels, defined as
$\omega_{mn}=\frac{1}{|X_{mn}|}\sum_{i=1}^{|X_{mn}|}(\hat{z}_{i}-z_{i})^{2},$ (4)
where $\hat{z}_{i}$ denotes the inferred label from the current global model and $z_{i}$ is its ground-truth label. If the value of $\omega_{mn}$ is large, it indicates that the current global model is not working well on local data or there are unseen samples. Therefore, updating the model by training on these data may lead to better performance of the global model. If the value of $\omega_{mn}$ is small, it means that the current data has little potential to further improve the global model since the knowledge of the data has already been learned or the data size is small. Note that to shorten the decision time, we use the potential value of the initial local data to be a proxy for the average potential value of all data within $T$. This is a reasonable approach when $T$ is not too long.
Figure 5: The AoI evolution of the models sent from MU $m$ to MSP $n$.
III-C2 Age of Information $\Delta$
The whole FL process involves model training, uploading, and waiting, which together constitute a complete cycle of model updating. From the perspective of MSP $n$, the age evolution of the model contributed by MU $m$ is shown in Fig. 5. Here, the filled circles denote the time instances when MU $m$ starts training the learning model. The intersections show the time instances when MSP $n$ receives the corresponding updated model, at which point the AoI drops to the lowest. We denote that at time $t$, the generation time of the last update received by MSP $n$ from MU $m$ is $s(t)$. Then, the instantaneous AoI of MU $m$’s learning model measured at MSP $n$, $\Delta_{mn}(t)$, is
$\Delta_{mn}(t)=t-s(t),$ (5)
which shows the elapsed time since the generation of the latest updated model. A smaller instantaneous AoI means that the latest model received is fresher at that time. To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models. With the aid of Figure 5, the average AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
[…]
$=\frac{1}{2}\tau_{n}+T_{mn}^{c}+T_{mn}^{t},$
where ${Q_{mn}}^{(r)}$ is the $r$-th trapezium under the curve. A lower average AoI indicates that the updated models are generally fresher over a long period.
Since MU $m$ constantly collects data throughout the FL process, the amount of training data collected for training in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time. The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computation resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
$T_{mn}^{c}=\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}.$ (7)
A smaller value of $\theta$ indicates higher accuracy, but leads to a higher MU cost, i.e., the number of local iterations, which is upper bounded by $\log(1/\theta_{m})$ [27].
Furthermore, the upload time $T_{mn}^{t}$ is directly related to the number of model parameters $b_{mn}$ and the resource $B_{mn}$ (i.e., uplink bandwidth) used to communicate between MU $m$ and MSP $n$, defined as
$T_{mn}^{t}=\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})},$ (8)
where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-interference-plus-noise ratio (SINR) for the communication between MU $m$ and MSP $n$. $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively. $\sigma^{2}$ is additive white Gaussian noise (AWGN).
Finally, by substituting (7) and (8) into (6), $\overline{\Delta}_{mn}$ can be expressed as
$\overline{\Delta}_{mn}=\frac{1}{2}\tau_{n}+\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}.$ (9)
IV Game Formulation
We first define the utility functions of MU $m$ and MSP $n$ in Section IV-A and Section IV-B, respectively. Then, we formulate the interactions among MUs and MSPs as an equilibrium problem with equilibrium constraints (EPEC) in Section IV-C, where the equilibrium criterion exists at both the level of MUs and MSPs due to the conflicting interests between them.
IV-A Utility of MU
We define the utility of MU $m$ composed of the rewards from MSPs and the cost incurred for contribution, which is also affected by resource constraints and immersion requirements for consuming vehicular metaverse services. The cost incurred due to local training and model uploading, which can be expressed as
$C_{mn}=c_{m}^{f}\log(1/\theta_{m})f_{mn}+c_{m}^{B}B_{mn},$ (10)
where $c_{m}^{f}$ and $c_{m}^{B}$ denote the cost factors of computation and communication resources, respectively. Let $\bm{f}_{m}\triangleq\left(f_{m1},f_{m2},\ldots,f_{mn}\right)$ and $\bm{B}_{m}\triangleq\left(B_{m1},B_{m2},\ldots,B_{mn}\right)$ be the computation and communication allocation profiles of MU $m$, respectively. Then, the utility maximization problem for MU $m$ within $T$ can be formulated as
Problem 1
$\max\,\,\Phi_{m}(\bm{f}_{m},\bm{B}_{m})=\sum_{n=1}^{N}(p_{mn}V_{mn}-C_{mn}),$ (11)
$s.t.$ $C1:\sum_{n=1}^{N}f_{mn}<f_{m}^{max},\,\,\sum_{n=1}^{N}B_{mn}<B_{m}^{max},$
$C2:\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}\leq\tau_{n},$
$C3:\frac{S_{m}}{f^{max}_{m}-\sum_{n=1}^{N}f_{mn}}<T^{req},$
where $S_{m}$ indicates the minimum computation resource required to enjoy other basic services. Constraints C1 and C3 indicate that the MU’s total computation and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services. Constraint C2 ensures that the total time for local training and uploading cannot exceed the time constraint $\tau_{n}$ for each round. The goal of each MU is to choose the optimal allocation of computation $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility. Note that $V_{mn}$ must be greater than 00. This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, then the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ will not participate in the trading.
IV-B Utility of MSP
Each MSP has a gain function $\psi$ associated with the IoM to measure the benefits of all the learning models from MUs to the service. In this paper, we adopt $\psi_{n}=\mu_{n}ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [28]. Let $\bm{p}_{n}\triangleq\left(p_{1n},\ldots,p_{mn}\right)$ and $\bm{p}_{\textrm{-}n}$ represent the reward profile of MSP $n$ and all other MSPs except MSP $n$, respectively. Then, the optimization problem for the MSP is defined as
Problem 2
$\max\,\,\Psi_{n}(\bm{p}_{n},\bm{p}_{\textrm{-}n})=\psi_{n}-\sum_{m=1}^{M}p_{mn}V_{mn},$ (12)
$s.t.\,\,p_{mn}>0,n\in\mathcal{N},m\in\mathcal{M},$
where $\mu_{n}$ refers to the profit conversion coefficient [29] from IoM, which is adjusted according to the AR services offered by the different MSPs. Note that the other MSPs’ decisions $\bm{p}_{\textrm{-}n}$ are captured by the MUs’ responses. Moreover, when MSP $n$ determine its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$). This thereby leads to reward competition among MSPs. Therefore, the optimization among MSPs can be considered as a non-cooperative game, termed a multi-MSP rewarding game $\Omega$ as follows.
Definition 1.
A multi-MSP rewarding game $\Omega$ is a tuple $\Omega=\left\{\mathcal{N},\bm{p},\bm{\Psi}\right\}$ defined by
  • Players: The set of MSPs;
  • Strategies: The reward decisions $\bm{p}_{n}$ of any MSP $n$;
  • Utilities: The vector $\bm{\Psi}=\left\{\Psi_{1},\Psi_{2},\ldots,\Psi_{n}\right\}$ contains the utility functions of all the MSPs defined in (12).
Figure 6: The structure of EPEC.
IV-C Multi-MSP Multi-MU Game as EPEC
In such a multi-MSP and multi-MU scenario, each MU has the choice of multiple MSPs to contribute learning models, and each MSP can incentivize multiple MUs to earn high-value learning models and achieve more benefits. Accordingly, each MSP competes with others for the equilibrium, which is constrained by the equilibrium among the MUs. This leads to the EPEC formulation [30].
As shown in Fig. 6, MUs and MSPs negotiate allocation strategies of computation and communication resources and reward decisions to maximize their benefits. In the upper level, MSPs determine the rewards they are willing to offer by considering their cost, the responses of MUs, and the decisions of other MSPs. In the lower level, each MU m gives the optimal allocation response for computation and communication resources by considering its resource constraints, cost, and rewards from different MSPs. The objective of EPEC is to find the equilibria at two levels, i.e., the point at which the MSPs’ (leaders’) utilities are maximized given that the MUs (followers) will choose their best responses. For the proposed EPEC, the equilibria at two levels are defined as follows.
Definition 2.
Let $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ denote the optimal computation and communication resource allocation of MU $m\in\mathcal{M}$ and the optimal reward decision of MSP $n\in\mathcal{N}$, respectively. Then, the points $(f^{*}_{mn},B^{*}_{mn})$ and $p^{*}_{mn}$ are the equilibria at two levels if the following conditions hold:
$\Phi_{m}((\bm{f}_{m}^{*},\bm{B}_{m}^{*}),\bm{p}_{n}^{*})\geq\Phi_{m}((\bm{f}_{m},\bm{B}_{m}),\bm{p}_{n}^{*}),\forall m\in\mathcal{M},$ (13)
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n}^{*},\bm{p}_{-n}^{*}\right)\geq$
$\Psi_{n}\left(\bm{f}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{B}_{m}^{*}(\bm{p}_{n}^{*},\bm{p}_{-n}^{*}),\bm{p}_{n},\bm{p}_{-n}^{*}\right),\forall n\in\mathcal{N}.$
In summary, the MSPs’ optimization problems are formulated as the following EPEC problems:
$\max_{p_{n}}\,\,\Psi_{n}=\psi_{n}-\sum_{m=1}^{M}\bigg{[}p_{mn}I_{mn}\bigg{(}\frac{1}{2}\tau_{n}$ (14)
$-\frac{x_{m}\tau_{n}\log(1/\theta_{m})}{f_{mn}^{*}}-\frac{b_{mn}}{B_{mn}^{*}\log_{2}(1+\varsigma_{mn})}\bigg{)}\bigg{]},$
[…]
To investigate the above EPEC, we address the lower level (Problem 1) and the upper level (Problem 2) by using the backward induction methods in the following section.
V Analysis and Solutions
In this section, we use backward induction to investigate the existence and uniqueness of the equilibria at two levels, i.e., the optimal computation and communication resource allocation of MUs at the lower level and the optimal reward decisions of MSPs at the upper level.
V-A Lower Level: Optimal Computation and Communication Resource Allocation Analysis
For any reward decisions $\bm{p}$ given by MSPs, MU $m$ would like to determine an optimal decision $\bm{f}_{m}^{*}$ and $\bm{B}_{m}^{*}$ to maximize its utility. We analyze and provide the solution for the optimal allocation of computation and communication resources as follows.
Definition 3.
If the problem is strictly concave for maximization and satisfies the following conditions, it will have a unique optimal solution [31]: (1)The Utility function is concave; (2) Constraints are all affine constraints; and (3) Strategy space is a compact and convex set.
Theorem 1.
Problem 1 has a unique optimal solution in the lower level, i.e., each MU has an optimal and unique allocation scheme for computation and communication resources.
Proof.
We first observe that the Hessian matrix of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $(f_{mn},B_{mn})$ is
$\bm{H}_{m}=\begin{bmatrix}\bm{H}_{m}^{f}&0\\ 0&\bm{H}_{m}^{B}\end{bmatrix},$ (15)
where $\bm{H}^{f}_{m}$ can be obtained by the second-order derivation of $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ with respect to $f_{mn}$
$\bm{H}^{f}_{m}$ $=\left[\frac{\partial^{2}\Phi_{m}(\bm{f}_{m},\bm{B}_{m})}{\partial f_{mn}\partial f_{mn^{\prime}}}\right]_{n,n^{\prime}\in\mathcal{N}}$ (16)
$=-diag\left(h^{f}_{m1},h^{f}_{m2}\ldots,h^{f}_{mn}\right)<\bm{0},$
where $h^{f}_{mn}=p_{mn}I_{mn}\frac{2x_{m}\tau_{n}}{f_{mn}^{3}}\log(1/\theta_{m})$.
Similarly, the $\bm{H}^{B}_{m}$ can be calculated as
$\bm{H}^{B}_{m}$ $=\left[\frac{\partial^{2}\Phi_{m}(\bm{f}_{m},\bm{B}_{m})}{\partial B_{mn}\partial B_{mn^{\prime}}}\right]_{n,n^{\prime}\in\mathcal{N}}$ (17)
$=-diag\left(h^{B}_{m1},h^{B}_{m2},\ldots,h^{B}_{mn}\right)<\bm{0},$
where $h^{B}_{mn}=\frac{2b_{mn}p_{mn}I_{mn}}{B_{mn}^{3}\log_{2}(1+\varsigma_{mn})}$.
We can easily show that $\bm{H}^{f}_{m}$ and $\bm{H}^{B}_{m}$ are negative semi-definite, and thus $\bm{H}_{m}$ is negative semi-definite and $\Phi_{m}(\bm{f}_{m},\bm{B}_{m})$ is concave. Moreover, according to the convexity condition of its constraints and Definition 3, it can be proved that Problem 1 has a unique optimal solution $(\bm{f}_{m}^{*},\bm{B}_{m}^{*})$. ∎
Next, we can obtain the best response function of MU $m$ by solving Problem 1. Based on the first-order condition and setting the derivative to zero, we can derive the following proposition.
Proposition 1. Given the digital currency price offered by the MSPs and the Lagrange multipliers induced by the constraints, the best response function $f_{mn}$ of MU $m$ satisfies:
$f_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},&\text{ $p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\,\,\text{otherwise},\end{cases}$ (18)
and the best response function $B_{mn}$ of MU $m$ satisfies:
$B_{mn}^{*}=\begin{cases}\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{$p_{mn}\geq\frac{F^{2}}{I_{mn}}$},\\ F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}},&\text{\,\,$0<p_{mn}<\frac{F^{2}}{I_{mn}}$},\\ 0,&\text{otherwise},\end{cases}$ (19)
where $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}$.
Proof.
First, we transform Problem 1 into an unconstrained optimization problem using the Lagrangian dual method. The Lagrangian function associated with MU $m$ is formulated as
$L_{m}\left(\bm{f}_{m},\bm{B}_{m},\bm{p}_{m},\lambda_{m},\beta_{m},\delta_{mn},\gamma_{m}\right)=\sum_{n}p_{mn}V_{mn}$ (20)
$-\sum_{n}\left[c_{m}^{f}f_{mn}log(1/\theta_{m})+c_{m}^{B}B_{mn}\right]$
$-\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})-\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})$
$-\sum_{n}\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]$
$-\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})],$
where $\lambda_{m}\geq 0,\beta_{m}\geq 0$, $\delta_{mn}\geq 0$ and $\gamma_{m}\geq 0$ are the Lagrangian dual variables corresponding to constraints C1, C2 and C3, respectively. Then, the first-order derivative of $L_{m}$ with respect to $f_{mn}$ and $B_{mn}$ can be derived as
$\frac{\partial L_{m}}{\partial f_{mn}}$ $=p_{mn}I_{mn}\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}-c_{m}^{f}log(1/\theta_{m})$ (21)
$-\lambda_{m}+\delta_{mn}log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-\gamma_{m}T^{req},$
and
$\frac{\partial L_{m}}{\partial B_{mn}}$ $=p_{mn}I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}-\beta_{m}$ (22)
$+\delta_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}.$
From (21) and the constraints in (11), the KKT conditions are given by
$\frac{\partial L_{m}}{\partial f_{mn}}=0,\,\,\lambda_{m}(\sum_{n}f_{mn}-f_{m}^{max})=0,$ (23)
$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
$\gamma_{m}[S_{m}-T^{req}(f_{m}^{max}-\sum_{n}f_{mn})]=0,$
$\lambda_{m},\delta_{mn},\gamma_{m}\geq 0,C1,C2,C3.$
Since $S_{m}<T^{req}\left(f_{m}^{max}-\sum_{n}f_{mn}\right)$ and $\sum_{n}f_{mn}<f_{m}^{max}$, we can obtain $\lambda_{m}=0$ and $\gamma_{m}=0$. Furthermore, $\frac{\partial L_{m}}{\partial f_{mn}}=0$ can be converted to
$(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-c_{m}^{f}=0.$ (24)
Similarly, the KKT conditions for $B_{mn}$ are obtained as
$\frac{\partial L_{m}}{\partial B_{mn}}=0,\,\,\beta_{m}(\sum_{n}B_{mn}-B_{m}^{max})=0,$ (25)
$\delta_{mn}[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}]=0,$
$\beta_{m},\delta_{mn}\geq 0,C1,C2.$
Since $\sum_{n}B_{mn}<B_{m}^{max}$, we can obtain $\beta_{m}=0$. Furthermore, $\frac{\partial L_{m}}{\partial B_{mn}}=0$ can be converted to
$(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-c_{m}^{B}=0.$ (26)
Then, based on the value of $\delta_{mn}$, the optimal allocation of computation and communication resources exists in the following two cases:
  • ($\delta_{mn}=0$): According to (24) and (26), we have
    $f_{mn}=\sqrt{\frac{p_{mn}I_{mn}x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=\sqrt{\frac{p_{mn}I_{mn}b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (27)
  • ($\delta_{mn}>0$): Substituting $f_{mn}$ and $B_{mn}$ into $\delta_{mn}\left[log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}log_{2}(1+\varsigma_{mn})}-\tau_{n}\right]=0$, we can get
    $\delta_{mn}=F^{2}-p_{mn}I_{mn},$ (28)
    $F=log(1/\theta_{m})\sqrt{\frac{x_{m}c_{m}^{f}}{\tau_{n}}}+\frac{\sqrt{b_{mn}c_{m}^{B}}}{\tau_{n}\sqrt{log_{2}(1+\varsigma_{mn})}}.$
    Next, bring $\delta_{mn}$ into $f_{mn}$ and $B_{mn}$, we have
    $f_{mn}=F\sqrt{\frac{x_{m}\tau_{n}}{c_{m}^{f}}},B_{mn}=F\sqrt{\frac{b_{mn}}{c_{m}^{B}log_{2}(1+\varsigma_{mn})}}.$ (29)
According to Proposition 1, the optimal strategy for each MU $m$ is influenced by five main factors: $p_{mn}$, $I_{mn}$, $c_{m}^{f}$, $c_{m}^{B}$ and $\tau_{n}$. Specifically, MU $m$ focuses on the rewards paid by MSPs, and the MU tends to put more computation and communication resources into the contribution when $p$ increases. Moreover, MUs are more willing to invest more resources when the potential contribution prediction $I_{mn}$ is higher, resulting in greater benefits. In addition, the virtual deadline $\tau_{n}$ set by the MSP $n$ determines the minimum criteria for resource allocation.
V-B Upper Level: Optimal Reward Decision Equilibrium Analysis
Given the responses $(\bm{f},\bm{B})$ from all MUs, and other MSPs’ decisions $\bm{p}_{\textrm{-}n}$, the optimal reward decision $\bm{p}_{n}$ of MSP $n$ can be obtained by solving Problem 2, defined as follows.
Proposition 2. Given $\bm{p}_{\textrm{-}n}$, MSP $n$’s best response strategy is
$\bm{p}_{n}^{*}=arg\max_{p_{mn}>0}\Psi_{n}\left(\bm{f}_{n}^{*}(\bm{p}_{n},\bm{p}_{-n}),\bm{B}_{n}^{*}(\bm{p}_{n},\bm{p}_{-n}),\bm{p}_{-n}\right).$ (30)
Then, we analyze the existence and uniqueness of the optimal reward decision equilibrium in Theorem 2.
Theorem 2.
There exists a unique Nash equilibrium (NE) [32] in the multi-MSP rewarding game $\Omega$.
Proof.
We define the Hessian matrix of $\Psi_{n}$ with respect to $\bm{p}_{n}$ as $(\bm{\Lambda}_{n}+\bm{H}_{n})$. The matrix $\bm{\Lambda}_{n}=diag\left(\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}^{2}},\ldots,\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}\right)$ and the second-order partial derivative matrix $\bm{H}_{n}$ is expressed by
$\bm{H}_{n}=\begin{bmatrix}0&\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}\partial p_{2n}}&\cdots&\frac{\partial^{2}\Psi_{n}}{\partial p_{1n}\partial p_{mn}}\\ \frac{\partial^{2}\Psi_{n}}{\partial p_{2n}\partial p_{1n}}&0&\cdots&\frac{\partial^{2}\Psi_{n}}{\partial p_{2n}\partial p_{mn}}\\ \vdots&\vdots&\ddots&\vdots\\ \frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{1n}}&\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{2n}}&\cdots&0\end{bmatrix},$ (31)
where
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$ $=\mu_{n}\frac{V_{mn}^{\prime\prime}(1+\sum_{m=1}^{M}V_{mn})-(V_{mn}^{\prime})^{2}}{(1+\sum_{m=1}^{M}V_{mn})^{2}}$ (32)
$-2V_{mn}^{\prime}-p_{mn}V_{mn}^{\prime\prime},\forall m\in\mathcal{M},$
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}=$ $-\mu_{n}\frac{V_{mn}^{\prime}V_{m^{\prime}n}^{\prime}}{(1+\sum_{m=1}^{M}V_{mn})^{2}},$ (33)
$\forall m\in\mathcal{M},m\neq m^{\prime}.$
To obtain the values of $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$, we need to obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$, $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$. From (24), we can obtain $\frac{\partial f_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}$, the steps of which are shown as follows. The first-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{x_{m}\tau_{n}}{f_{mn}^{2}}-2(p_{mn}I_{mn}+\delta_{mn})\frac{x_{m}\tau_{n}}{f_{mn}^{3}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}=0$. Consequently, we can obtain
$\frac{\partial f_{mn}}{\partial p_{mn}}=\frac{I_{mn}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (34)
The second-order derivative of $\frac{\partial L_{m}}{\partial f_{mn}}=0$ with respect to $p_{mn}$ is expressed as
$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial f_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}f_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (35)
by substituting (34) into (35), we have
$\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}f_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (36)
Likewise, we derive $\frac{\partial B_{mn}}{\partial p_{mn}}$ and $\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}$ with the following steps based on (26). The first-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ is expressed as $I_{mn}\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}-2(p_{mn}I_{mn}+\delta_{mn})\frac{b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}=0$. Accordingly, we have
$\frac{\partial B_{mn}}{\partial p_{mn}}=\frac{I_{mn}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})}>0.$ (37)
Similarly, we obtain the second-order derivative of $\frac{\partial L_{m}}{\partial B_{mn}}=0$ with respect to $p_{mn}$ as
$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{I_{mn}\frac{\partial B_{mn}}{\partial p_{mn}}(p_{mn}I_{mn}+\delta_{mn})-I_{mn}^{2}B_{mn}}{2(p_{mn}I_{mn}+\delta_{mn})^{2}},$ (38)
by substituting (37) into (38), we have
$\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}=\frac{-I_{mn}^{2}B_{mn}}{4(p_{mn}I_{mn}+\delta_{mn})^{2}}<0.$ (39)
Based on (34), (36), (37), and (39), we can easily infer that $V_{mn}^{\prime}>0$ and $V_{mn}^{\prime\prime}<0$, as shown in (40). Then, we can obtain $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}\partial p_{m^{\prime}n}}<0$. For $\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}$, We substitute the specific expressions for $V_{mn}^{\prime}$ and $V_{mn}^{\prime\prime}$ into (32) and obtain the following result as shown in (41). Next, we randomly choose a vector $\bm{h}\in\mathbb{R}^{M\times 1}$ with elements not all 00. For $\bm{\Lambda}_{n}$, it is easy to see that $\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}=\sum_{m}\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}(h^{m})^{2}<0$. Furthermore, $\bm{h}^{T}\bm{H}_{n}\bm{h}=\frac{-\mu_{n}}{(1+\sum_{m}V_{mn})^{2}}\left(\sum_{m}h^{m}V_{mn}^{\prime}\right)^{2}<0$. Thus, we have $\bm{h}^{T}(\bm{\Lambda}_{n}+\bm{H_{n}})\bm{h}=\bm{h}^{T}\bm{\Lambda}_{n}\bm{h}+\bm{h}^{T}\bm{H}_{n}\bm{h}<0$, indicating that the utility function $\Psi_{n}$ is strictly concave. So far, the uniqueness of the Nash equilibrium solution to the multi-MSP rewarding game $\Omega$ has been proved.
$V_{mn}^{\prime}=\frac{\partial V_{mn}}{\partial p_{mn}}=I_{mn}\left(\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial f_{mn}}{\partial p_{mn}}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial B_{mn}}{\partial p_{mn}}\right)>0,$ (40)
$V_{mn}^{\prime\prime}=\frac{\partial^{2}V_{mn}}{\partial p_{mn}^{2}}=I_{mn}\left[\frac{-2x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{3}}\cdot(\frac{\partial f_{mn}}{\partial p_{mn}})^{2}+\frac{x_{m}\tau_{n}log(1/\theta_{m})}{f_{mn}^{2}}\cdot\frac{\partial^{2}f_{mn}}{\partial p_{mn}^{2}}\right.$
$\left.-\frac{2b_{mn}}{B_{mn}^{3}log_{2}(1+\varsigma_{mn})}\cdot(\frac{\partial B_{mn}}{\partial p_{mn}})^{2}+\frac{b_{mn}}{B_{mn}^{2}log_{2}(1+\varsigma_{mn})}\cdot\frac{\partial^{2}B_{mn}}{\partial p_{mn}^{2}}\right]<0.$
$\frac{\partial^{2}\Psi_{n}}{\partial p_{mn}^{2}}=-\left[\frac{4a_{1}I_{mn}^{2}\delta_{mn}+a_{1}I_{mn}^{3}p_{mn}}{4f_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+\frac{4b_{mn}I_{mn}^{2}\delta_{mn}+b_{mn}I_{mn}^{3}p_{mn}}{4a_{2}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}}+I_{mn}^{2}\frac{\left(\frac{a_{1}a_{2}I_{mn}B_{mn}+b_{mn}I_{mn}f_{mn}}{2a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})}\right)^{2}}{(1+\sum_{m}V_{mn})^{2}}\right.$ (41)
$\left.+\frac{3I_{mn}^{3}(a_{1}a_{2}B_{mn}+b_{mn}f_{mn})}{4a_{2}f_{mn}B_{mn}(\delta_{mn}+I_{mn}p_{mn})^{2}(1+\sum_{m}V_{mn})}\right]<0,\text{where}\,a_{1}=x_{m}\tau_{n}log(1/\theta_{m}),a_{2}=log_{2}(1+\varsigma_{mn}).$
1
2Input: Action threshold $[p_{min},p_{max}]$, learning rate;
3 Output: The reward decisions $\bm{p}$ of MSPs, resource allocation $[\bm{f},\bm{B}]$ of MUs;
4 Initialization: local observation $s_{n}$, actor network $\bm{\alpha}_{n}$, critic network $\bm{\beta}_{n}$ and episode buffer $\mathcal{D}_{n}$ for each MSP agent $n$ ;
5 # Trading process using MDDR:
6 for  Episode = $1,2,\ldots$ do
7      for Epoch $k=1,\ldots,|\mathcal{D}|$ do
8            foreach MSP agent $n\in\mathcal{N}$ do
9                  Observe space $s_{n}(k)$;
10                   Choose price $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$ by sampling from $\pi_{\alpha_{n}}(s_{n}(k))$ ;
11             end foreach
12            All MSP agents do:
13             Perform a joint reward $\bm{p}(k)=\left\{\bm{p}_{1}(k),\ldots,\bm{p}_{n}(k)\right\}\in\mathcal{A}$;
14             Broadcast reward decisions;
15             Interact with the environment and receive responses from the MUs ;
16             Calculate utility according to (12) ;
17             Store $s_{n}(k)=[V_{n}(k),\bm{p}_{n}(k),\mathcal{U}_{n}(k),V_{n}(k+1)]$ ;
18       end for
19      Update actor network $\bm{\alpha}$ and critic network $\bm{\beta}$ for each MSP agent as [33] ;
20       Clear up relay buffer $\mathcal{D}$ ;
21 end for
22# FL training guided by trading results within $T$:
23 foreach MSP agent $n\in\mathcal{N}$  do
24      foreach MU $m\in\mathcal{M}$ do
25             Based on the trading results, utilize $f_{mn}$ and $B_{mn}$ to locally train the task of MSP $n$, and subsequently upload the model to the MSP $n$.
26       end foreach
27      Aggregate and issue an updated global model;
28       Supply the updated model to the AR engine;
29 end foreach
Algorithm 1 Entire Process of the Framework.
We verify the existence and uniqueness of the equilibrium solution for EPEC through Theorem 1 and Theorem 2. However, balancing multiple and conflicting interests, i.e., the utilities of MSPs and MUs, inevitably complicates the optimization problem. To find a solution that optimizes the profits of MSPs while taking into account the utilities of MUs, we develop a reinforcement learning-based strategy in the following section.
VI Dynamic Reward Strategy for MSPs
According to the EPEC proved above, it is possible to obtain an optimal solution. However, the optimal approach faces the following practical difficulties: (i) under the time-varying network, the optimal approach is time-consuming due to the complexity of (12), which is non-linear with complicated form and the reward decisions of MSPs are tightly coupled. (ii) MSPs must know private information about MUs (e.g., cost factors) as a prior condition for determining their rewards, which violates the privacy of MUs. To overcome the above difficulties, we further formulate the multi-MSP rewarding game as a multi-agent Markov decision process (MAMDP) [34] to adapt to dynamic channel conditions, where each MSP acts as an individual agent to make intelligent reward decisions.
VI-A MAMDP for Multi-MSP Rewarding Game $\Omega$
We train the reward model based on the state-of-the-art policy gradient method proximal policy optimization (PPO) for the reasons described in [35]. The $MAMDP=\left\langle\mathcal{S}_{n},\mathcal{A}_{n},\mathcal{P}_{n},\mathcal{U}_{n}\right\rangle$ for multi-MSP rewarding game is composed of state space $\mathcal{S}_{n}\triangleq\left\{s_{n}\right\}$, action space $\mathcal{A}_{n}\triangleq\left\{\bm{p}_{n}\right\}$, state transition probability $\mathcal{P}_{n}\triangleq\left\{P_{n}\right\}$, and utility $\mathcal{U}_{n}\triangleq\left\{\Psi_{n}\right\}$. For the MSP agent, the local observation contains the channel information and the responses of MUs, all of which are captured in the IoM of (2). Thus, we set the local observation of MSP $n$ at the $k$-th stage game defined as $s_{n}(k)=[V_{1n}(k),V_{2n}(k),\ldots,V_{mn}(k)]$, and the state of the environment is $\mathcal{S}(k)=[s_{1}(k),\ldots,s_{n}(k)]$. At the $k$-th stage game, MSP $n$ observes a state $s_{n}(k)$ and determines an action $\bm{p}_{n}(k)$ within $[p_{min},p_{max}]$. When an action $\bm{p}_{n}(k)$ is applied to state $s_{n}(k)$, the agent $n$ receives a utility $\mathcal{U}_{n}(k)$ from the environment. Taking into account the competition among agents, we define the reward $\mathcal{U}_{n}$ for each MSP by (12). The state transition probability $P(s_{n}(k+1)\mid s_{n}(k),\bm{p}_{n}(k))$ leads to the new state $s_{n}(k+1)$ after executing an action $\bm{p}_{n}(k)$ at the state $s_{n}(k)$.
Figure 7: The utilities of MSPs.
(a) The utilities of MUs
(b) The reward decisions of MSP 1
(c) The reward decisions of MSP 2
(d) The reward decisions of MSP 3
Figure 8: The convergence processes of MDDR under dynamic networks.
VI-B The Multi-agent DRL-based Dynamic Reward (MDDR)
In our work, each MSP agent operates the DRL-based dynamic reward algorithm in a fully distributed manner. Algorithm 1 shows the pseudocode for model trading using MDDR under a dynamic environment (Lines 5-19) and the guidance for the FL process (Lines 22-27).
Trading process: When the game starts, each MSP $n$ initializes its actor $\bm{\alpha}_{n}$ and critic network $\bm{\beta}_{n}$, essential parameters and inference loss for trading (Line 3). During each game, each agent feeds the observed information into the policy network $\bm{\alpha}_{n}$ to derive its reward policy (Lines 8-9). Then, all the agents generate reward decisions and broadcast them to the MUs (Lines 12-13). After completing these steps, the MSP agents enter the interaction phase with the environment, receive the best feedbacks from the MUs, calculate their utilities, and store information into episode buffer $\mathcal{D}$ (Lines 14-16). The experience is collected through the above interactive process (Lines 8-16) until the buffer $\mathcal{D}$ is full. Then, the critic-network $\bm{\beta}_{n}$ evaluates the actor-network $\bm{\alpha}_{n}$ based on the experience and in turn adjusts it (Line 18). The game will terminate when the maximum number of training episodes is reached and then the trading results will be obtained.
FL training guided by trading results: After obtaining the trading results (MSPs’ reward decisions $\bm{p}$ and resource allocation $(\bm{f},\bm{B})$), MUs and MSPs enter the FL phase. MUs follow the resource allocation scheme for local training and model uploading (Line 24). Subsequently, the MSP is responsible for aggregating and updating the global model (Line 26). Finally, the updated model is provided for use by the AR engine (Line 27). This process iterates multiple times until the total FL time $T$ is reached. During real-time interactions, the AR service of the MSP can be enhanced in terms of immersion and service quality by the updated model.
VI-C The Complexity and Convergence of MDDR
The complexity of MDDR is mainly influenced by the structure of the actor-critic network. According to [36], the total computational complexity of the fully connected layer can be expressed as the number of multiplications $\mathcal{O}(\sum_{u=1}^{U}\varrho_{u}\cdot\varrho_{u-1})$, where $\varrho_{u}$ is the number of neurons in the $u$-th hidden layer. In this work, each agent has its actor-critic network, structured as $[5,256,10]$ and $[5,256,1]$, respectively. Therefore, its complexity can be expressed as $\mathcal{O}(5\times 256+256\times 10+5\times 256+256\times 1)$ without burden on the MSP. After several attempts, the MDDR can achieve convergence [37] (as per Fig. 7 and Table II) with a suitable set of hyper-parameters.
TABLE II: The Utilities of three reward approaches
MSP 1 MSP 2 MSP 3 Total
Optimal 2169.99 2564.56 3035.96 7770.51
MDDR 2169.11 2563.95 3034.94 7768.00
MAPPO 2169.23 2563.96 3035.26 7768.45
(a) MSP1
(b) MSP2
(c) MSP3
(d) The IoM for MSP 1
(e) The IoM for MSP 2
(f) The IoM for MSP 3
Figure 9: Evaluation results for MNIST.
(a) MSP1
(b) MSP2
(c) MSP3
(d) The IoM for MSP 1
(e) The IoM for MSP 2
(f) The IoM for MSP 3
Figure 10: Evaluation results for GTSRB.
VII Performance Evaluation
First, we verify the near-optimal performance of MDDR. Then, to evaluate the effectiveness of the proposed framework, we compare its performance with benchmark schemes in object detection and classification for AR services through FL.
VII-A Simulation Settings
[…]
VII-B Benchmark Schemes
To our knowledge, there are no comparable solutions to the model trading problem in the vehicular metaverse. Therefore, under the uniform rewards of MSPs, we develop and extend four benchmark schemes for multi-MSP and multi-MU scenarios while meeting FL deadlines and resource constraints of devices.
  • x_based: Inspired by the incentive mechanism of the Stackelberg game for federated learning [25], the MU allocates computation and communication resources based on the amount of data involved in local training.
  • w_based: The MU adjusts the resource allocation according to the inference loss [42], i.e., the potential value of local data in our work.
  • w_x_based: Combining the x_based and w_based designs, the MU allocates resources considering both the data size and the potential value of data.
  • fixed: Taking into account the differences in performance of the various devices that exist in reality, we randomly allocate a fixed set of computation and communication resources to reflect the heterogeneity of devices.
TABLE III: Time comparison of five schemes for achieving the same performance.
Schemes MNIST GTSRB
90% Accuracy 50% mAP 70% mAP50
MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3 MSP 1 MSP 2 MSP 3
immersion-aware 15.23s 13.09s 7.59s 513.16s 581.09s 623.89s 418.28s 447.69s 451.27s
x_based 17.79s 18.03s 19.80s 812.53s 898.59s 1061.50s 596.57s 679.91s 890.47s
w_based 30s+ 24.04s 10.28s 1064.83s 1175.58s 1200s+ 774.09s 953.50s 1013.89s
w_x_based 23.20s 19.08s 12.34s 901.61s 1190s 1194s 596.35s 848.68s 932.73s
fixed 30s+ 30s+ 30s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+ 1200s+
  • The $1200$s+ and $30$s+ indicate that the guidance time for the current trading is exceeded.
VII-C Performance Evaluation of MDDR
First, we consider a relatively static communication environment (fixed channel conditions) to show the effectiveness of MDDR compared to the optimal solution and the general multi-agent depth approximation policy optimization (MAPPO) approach. We trained our proposed MDDR and MAPPO until convergence, and the results are shown in Fig. 7 and Table II. We find that both MDDR and MAPPO achieve good results, but MAPPO slightly outperforms MDDR. This is because MAPPO is a centrally controlled distributed execution [33] that can utilize more information for decision making. However, this approach is not suitable for non-cooperation due to the need to share parameters among MSPs. For the fully distributed MDDR, it is more suitable for non-cooperative scenarios. The reason is that in MDDR, each agent has a self-contained actor-critic network, and information is not required to be shared. This suggests that the approach can achieve considerable utility while preserving privacy and saving communication costs significantly. Then, we deploy the MDDR in dynamic communication environments where channel conditions change over time. The behaviors of each MU and MSP in the trading are shown in Figs. 8(a), 8(b), 8(c) and 8(d), reflecting that the MDDR can converge rapidly and show good stability in dynamic network environments.
VII-D Benchmark Comparison
Figs. 9(a), 9(b), and 9(c) show the performance improvement in classification accuracy over time, where the immersion-aware outperforms the other schemes. This is because immersion-aware utilizes IoM to incentivize MUs to allocate computation and communication resources for providing learning models to MSPs. The IoM metric not only evaluates the freshness and accuracy of the model but also considers the amount and potential value of raw data used for training. As a result, resource-limited MUs are motivated to contribute more resources, providing valuable learning models to MSPs at an accelerated pace. In contrast, baseline schemes like x_based, w_based and w_x_based struggle to offer adequate incentives for MUs to contribute their available resources due to uniform rewards. For example, if a given MU has already reached the maximum resource it can contribute when the reward is $p$, then even if $p$ is increased further, the MSP’s gains will be impaired rather than increased. Consequently, the uniform reward $p$ will not be increased, leading to a situation where other MUs with available resources will not contribute more resources for providing higher-value learning models.
In addition, Figs. 9(a), 9(b), and 9(c) reveal that MSP 3 performs better than MSP 1 and MSP 2. For example, at the $10$th second, the accuracy of MSP 3 is about $88\%$, which is $14\%$ and $3\%$ higher than those of MSP 1 and MSP 2, respectively. This is because the potential value $w$ is higher for MSP 3, leading to more rewards being given by MSP 3. As a result, the MUs will allocate more computation and communication resources to provide learning models for MSP 3 than for MSP 1 and MSP 2. For the three different MSPs, some fluctuations are normal for model training (e.g., at time $10$s). It can be seen that the effect of fixed is the worst due to synchronous FL, which is limited by the MU with the least resource. Meanwhile, it is clear from Figs. 9(d)9(e) and 9(f) that the nature of the effectiveness of the five schemes is influenced by IoM. With the higher IoM, the performance is improved faster, resulting in a better immersive experience. Yet, the IoM of w_based in Fig. 9(e) is higher than x_based and w_x_based but the performance improvement rate is not as fast as them. The reason is that the IoM here represents the sum of all IoM from MUs, whereas the rate of performance improvement is hampered by the worst MU. Additionally,from Figs. 9(b) and 9(c), it can be observed that both w_x_based and w_based perform better when compared to Fig. 9(a). The main reason is the higher potential value $w$ associated with MSP 2 and MSP 3.
For object detection on GTSRB, Figs. 10(a), 10(b) and 10(c) show the mean average precision (mAP) and Table III gives the mAP50 (The value of mAP when the intersection over union (IOU) threshold is larger than $0.5$) results. We observe that compared with MNIST, the performance enhancement for object detection is slower due to the complexity of the detection model. Specifically, it takes at least $513.16$ seconds to achieve a $50\%$ mAP using the immersion-aware method. Our research introduces the flexibility to adjust decision guidance time, making it adaptable to various task types. Meanwhile, Figs. 10(d)10(e)10(f) and Table III demonstrate that MSP 1 with greater IoM takes less time to achieve the same mAP, which indicates that IoM works well to capture immersion. For instance, MSP 1 has the highest IoM, and the time that it takes to reach $50\%$ mAP is $513.16$s, which is $67.93$s and $110.73$s less than those of MSP 2 and MSP 3, respectively. The above comparison validates the feasibility of the proposed trading framework and the advantages of IoM-based metric incentives.
VIII Conclusion
To facilitate the development of vehicular metaverse services, we propose an immersion-aware model trading framework that incorporates FL to incentivize MUs to become contributors and provide locally trained data for AR services, i.e., object detection and classification. Specifically, we construct an EPEC problem with MSPs as leaders and MUs as followers to achieve an equilibrium among their interests. A new metric called “IoM” is designed to comprehensively evaluate the enhancement brought by the learning models of MUs for AR services. In addition, considering the competitive relationship among MSPs and the dynamic network environment, we develop a fully distributed MDDR approach to obtain the reward decisions of MSPs. Through extensive simulations using AR-related vehicle and MNIST datasets, our results demonstrate that the proposed framework can efficiently provide MUs with better immersive experiences for AR services in the vehicular metaverse.